OpenStudy (anonymous):

Find the radius of the largest circle with centre (1,0) that can be inscribed in the ellipse \[x ^{2}/16 + y ^{2}/4 = 1\]

OpenStudy (anonymous):

2? If I'm imagining it right

OpenStudy (anonymous):

sorry, @Ishaan, thats not the answer

OpenStudy (anonymous):

Oh, Circle's center (1,0) I took it as (0,0)

OpenStudy (anonymous):

\[(x-1)^2 + y^2 = r^2\]\[x^2 -2x + 1 +y^2 = r^2\]\[x^2 - 2x + y^2 =r^2 -1\]Hmm how about sqrt(5)?

OpenStudy (anonymous):

answer: \[\sqrt{11}/3\] wanna know the process..

OpenStudy (anonymous):

No :/ I will have to solve it now. I got the answer but had to use Calculus, do you know of a better way, something graphical?

OpenStudy (anonymous):

or Geometrical?

OpenStudy (experimentx):

well, i think it can be done this way, the maximum radius of the circle will the the smallest distance from the center of circle to the curve. now our point is (1,0), let x, y be the point in the ellipse. the distance is d = sqrt{(x-1)^2 + y^2} --> if we change this y into x's then we will have a function a function in x, minimizing the value taking extremum might give us this point x, and so we might find corresponding y, hence the radius of the circle.

OpenStudy (anonymous):

I did the same

OpenStudy (experimentx):

best of luck isshan ... i hope you will share the result with me.

OpenStudy (anonymous):

Thanks experimentx! I did fine today.